OpenStudy (anonymous):

A stadium has 48,000 seats. Seats sell for $35 in Section A, $20 in Section B, and $15 in Section C. The number of seats in Section A equals the total number of seats in Sections B and C. Suppose the stadium takes in $1,273,000 from each sol-out event. How many seats are in each section?

OpenStudy (anonymous):

I will fan and medal for anyone that can help.

OpenStudy (anonymous):

Okay. Algebra: Linear Algebra (NOT Linear Equations) Solvers Lessons Answers archive Click here to see ALL problems on Linear Algebra Question 535536: A stadium has 49000 seats. Seats sell for $25 in Section A, $20 in Section B, and $15 in Section C. The number of seats in Section A equals the total number of seats in Section B and C. Suppose the stadium takes in $1052000 from each sold-out event. How many seats does each section hold? (Scroll Down for Answer!) Did you know that Algebra.Com has hundreds of free volunteer tutors who help people with math homework? Anyone can ask a math question, and most questions get answers! OR get immediate PAID help on: Go!!! Answer by mananth(14698) About Me (Show Source): You can put this solution on YOUR website! A stadium has 49000 seats. Seats sell for $25 in Section A, z $20 in Section B,-------------------x seats $15 in Section C. ------------------y seats (x+y)=z 25(x+y)+20x+15y=1052000 25x+25y+20x+15y=1052000 45x+40y=1052000 /5 9x+8y=210400------------------1 2x+2y=49000 /2 x+y=24500 multiply by -9 -9x-9y=-220500----------------2 add equation (1) & (2) -y=-10100 y= 10100 Section C plug value of y in equation x+y = 24500 x+10100=24500 x=14400 seats Section B x+y = Z= 24500 Section A CHECK 24500+14400+10100= 49,000 24500*25+14400*20+15*10100=1,052,000

OpenStudy (anonymous):

Hope this helped! Have a great day!